OpenStudy (precal):

Relative minimum f(theta)=square root 3 times theta - 2 sin(theta)

OpenStudy (precal):

\[f (\theta)=\sqrt{3} \theta-2\sin \theta\]

OpenStudy (dan815):

take derivative wrt to theta

OpenStudy (precal):

I found the derivative and set it to zero \[f ' (\theta)=\sqrt{3}-2\cos(\theta)\]

OpenStudy (precal):

I thought the answer was pi/6

OpenStudy (dan815):